OpenStudy (akohl):

Two cellos are being played in a duet. At a certain point, one cello plays a note of frequency 196 Hz and the other plays a note of frequency 220 Hz. What is the period of the beat? 0.1 s 0.4 s 0.25 s 4 s

OpenStudy (akohl):

@aaronq I keep getting .04s

OpenStudy (aaronq):

the frequency is 4, right? \(\sf T=\dfrac{1}{f}=\dfrac{1}{4/s}=0.25~s\)

OpenStudy (akohl):

isn't the frequency 24

OpenStudy (aaronq):

oh yeah, i read "200 Hz" not 220, my bad. yeah it's 0.041666 s which is none of the options.

OpenStudy (akohl):

K